OpenStudy (anonymous):

Find the Derivative:

OpenStudy (anonymous):

\[f(x)=(\frac{ x-5 }{ x ^{2}+2x})^{2011}\]

OpenStudy (anonymous):

what I got was \[2011(\frac{ x-5 }{ x ^{2} +2x})^{2010} * (\frac{ 1 }{ 2x+2 })\]

OpenStudy (anonymous):

need the quotient rule for the derivative of \[\frac{x-5}{x^2+2x}\]

OpenStudy (anonymous):

you can't take the derivative of a quotient by taking the derivative of the numerator and denominator separately